# Inferring CSM Properties of Type II SNe Using a Magnitude-Limited ZTF Sample

### v2

@@ Abstract
supernovae, mass-loss,
@@ 1 Introduction
Light curves of core-collapse supernovae (CCSNe), Type II SNe in particular, exhibit a large amount of diversity, varying across orders of magnitude in rise times, luminosities, and durations. The simple progenitor scenario, in which the initial mass is the only factor affecting the SN type or its light curve, cannot adequately explain the extensive observational diversity we see in photometry and spectroscopy – particularly with the acknowledgement of the role binarity plays in stellar evolution (e.g. Eldridge et al., 2008; Sana et al., 2012; Eldridge et al., 2018; Zapartas et al., 2019; Zapartas et al., 2021) via binary induced mass transfer and mergers.
An area being explored in greater detail is the degree to which diversity arises from stars with similar initial masses and evolutionary histories that, nonetheless, produce distinct observational signatures at the time of explosion; e.g. varying mass of H envelopes, progenitor radii and H-richness of the outer envelope (e.g. Popov, 1993; Chieffi et al., 2003; Young, 2004; Reynolds et al., 2020; Humphreys et al., 2020; Hiramatsu et al., 2021; Moriya et al., 2023; Dessart & Jacobson-Galán, 2023). Type II SNe result from the core-collapse of stars with initial masses between 8–20 $\textup{M}_{\odot}$ (e.g. Eldridge & Tout, 2004; Smartt, 2009, 2015; Van Dyk, 2017; Beasor et al., 2020). The most common subtype, Type IIP, originate from red supergiants (RSGs) – a connection confirmed through pre-explosion HST imaging (see Smartt, 2009, 2015). Their light curves exhibit H-recombination powered $\sim$100 d plateaus following steep rises to peak brightness, typically occurring within days to a week (e.g. Langer, 2012; Anderson et al., 2014; González-Gaitán et al., 2015; Rubin et al., 2016; Valenti et al., 2016).
@@
From the emergence of narrow emission lines in early spectra of young SNe (flash ionisation; Gal-Yam et al., 2014; Groh, 2014; Yaron et al., 2017; Gal-Yam, 2017; Bruch et al., 2023), strong evidence has been presented for the presence of a substantial mass of dense material close to the progenitor at the time of core-collapse. Narrow lines are likely the result of shock breakout (SBO) shock-heating and ionising a slow-moving, dense material (e.g. Yaron et al., 2017; Irani et al., 2024; Jacobson-Galán et al., 2024a). As the narrow lines typically persist for only a $\sim$ few days post-explosion, it is assumed that the CSM required is the result of mass-loss from the star in the years immediately preceding core-collapse (e.g. Das & Ray, 2017; Davies et al., 2022; Tinyanont et al., 2022; Pearson et al., 2023).
@@
Measurements of the CSM properties from the flash ionisation allow for constraints on the mass-loss rate, $\dot{M}$, and late-stage RSG instabilities experienced in the centuries-decades-years immediately before core-collapse (e.g. Mauron & Josselin, 2011; Yaron et al., 2017; Morozova et al., 2018; Stroh et al., 2021; Pearson et al., 2023; Tinyanont et al., 2022; Bruch et al., 2021, 2023; Moriya et al., 2023). Bruch et al. (2021, 2023) find that $\sim$ 60% of Type II SNe show evidence for significant amounts of dense CSM confined to a region around the progenitor at the time of explosion – this figure is not corrected for observational biases and not volume limited. Potential precursor events (e.g. Fraser et al., 2013; Jacobson-Galán et al., 2022; Dong et al., 2024; Warwick et al., 2025) provide further evidence of eruptions close to the ‘classical’ core-collapse event.
@@
This high fraction of progenitors with ‘enhanced’ $\dot{M}$ is supported by detailed studies of nearby events: SN 2021yja ($\approx$23 Mpc; Kozyreva et al., 2022; Hosseinzadeh et al., 2022), SN 2023ixf ($\approx$7 Mpc; Zimmerman et al., 2024; Hosseinzadeh et al., 2023; Hiramatsu et al., 2023; Jencson et al., 2023; Jacobson-Galán et al., 2023; Li et al., 2024; Singh et al., 2024) and SN 2024ggi ($\approx$7 Mpc; Chen et al., 2024b; Xiang et al., 2024; Pessi et al., 2024b; Jacobson-Galán et al., 2024b), which, combined with early photometric and spectroscopic data, confirm CSM around their RSG progenitors. In these cases, dense, optically thick CSM causes the SBO to occur within the CSM rather than at the stellar surface (Förster et al., 2018; Tinyanont et al., 2022; Pearson et al., 2023), producing rapid rises and enhanced peak luminosities (e.g. Moriya et al., 2011; Das & Ray, 2017; Morozova et al., 2017, 2018; Bruch et al., 2021, 2023; Tinyanont et al., 2022; Pearson et al., 2023; Moriya et al., 2023; Li et al., 2024). For slower-rising events ($\geq$ 5 d), the progenitor radius also significantly influences the rise time by affecting the SBO time and location (González-Gaitán et al., 2015; Morozova et al., 2016; Moriya et al., 2023; Irani et al., 2024).
@@
$\dot{M}$ for RSGs have been measured through multiple techniques: mid-IR observations of circumstellar dust in clusters show $\dot{M}$ $\sim$10-6 – 10-5 $\textup{M}_{\odot}$ yr-1 (e.g. Beasor & Davies, 2018; Beasor et al., 2020), consistent with rates derived from molecular line and radio measurements (e.g. Mauron & Josselin, 2011) and comparing pre-explosion progenitor properties to theoretical stellar evolution models (e.g. Smartt, 2009). Type IIn progenitors exhibit much higher rates of 10-3 – 1 $\textup{M}_{\odot}$ yr-1, derived from multi-wavelength observations (e.g. Kiewe et al., 2012; Taddia et al., 2013; Fransson et al., 2014), and combined X-ray, radio, and spectroscopic signatures (e.g. Smith, 2017a, b).
@@
However, $\dot{M}$ inferred from RSG observations alone are insufficient to produce the measured $M_{CSM}$ and $R_{CSM}$ on the timescale of decades to months pre-explosion (e.g. Bruch et al., 2021, 2023). Popular mechanisms for end-of-life mass-loss include: wave-driven energy heating into the stellar envelope (e.g. Fuller, 2017; Morozova et al., 2020; Wu & Fuller, 2021), radiation-driven mass-loss (e.g. Vink, 2008; Vink, Jorick S. & Sabhahit, Gautham N., 2023), instabilities caused by explosive shell burning (e.g. Arnett & Meakin, 2011; Smith & Arnett, 2014), common envelope interactions caused by binary interactions (e.g. Chevalier, 2012; Sana et al., 2012) and RSG ‘superwinds’ (e.g. Davies et al., 2022).
@@
Previous studies have been successful in characterising smaller subsets of Type II SN properties such as luminosities and rise times (e.g. Taddia et al., 2013; Anderson et al., 2014; Sanders et al., 2015; Gall et al., 2015; Rubin et al., 2016; Valenti et al., 2016; Graur et al., 2017; Davis et al., 2019) but are limited to small numbers or incomplete samples, made up of well-observed SNe detected in heterogeneous galaxy-targeted surveys. Infrequent and inconsistent survey cadence lead to inadequate coverage on the rise, limiting the amount of information one can infer from the rising light curve. The wide area, high cadences and untargeted nature of modern surveys allow for larger, more complete samples to be curated – lending to more detailed statistical analysis of population characteristics and their frequencies. This work builds upon previous studies by making use of robust statistical methods and large, highly complete surveys.
@@ 2.3 Gaussian Process Regression
CCSN light curves are difficult to model due to the extensive variety in their photometric behaviour. There exist analytical attempts to address the problem using parametric fitting functions (e.g. Villar et al., 2017; Villar et al., 2019) and there are advancements in theoretical models to produce synthetic light curves with more likeness to observed light curves (Morozova et al., 2017; Das & Ray, 2017; Moriya et al., 2023, and references therein). Whilst certain parameterisation and generalised empirical models have proven to be adequate in some scenarios, few of these models can fully characterise the diversity of parameters and properties present in transients being uncovered by large surveys. Thus, we are motivated to use a non-parametric technique such as Gaussian process regression (GPR; Rasmussen et al. 2004).
GPR is a non-parametric, Bayesian machine learning method for modelling data with functions of an unknown form (see Aigrain & Foreman-Mackey, 2023, for a review). For single-band SN light curve interpolation, the unknown function a 1-dimensional Gaussian process (GP) approximates is flux as a function of time. We include the effective wavelength, $\lambda_{\rm{eff}}$, of each filter band, and train in 2-dimensions – e.g. flux as a function of time and effective wavelength – which is often expressed as probability by Eq. 1, similar to methodology used in Thornton et al. (2024).
@@
For modelling diverse SN light curves, we use a Matèrn-5/2 covariance function (K) that captures both smooth evolution in addition to the sharp transitions characteristic of SNe – the kernel includes an additive white noise term to account for photometric uncertainties . We implement the GPR using the python package george (Ambikasaran et al., 2015), combining the Matèrn-5/2 kernel with the function from Villar et al. (2019) – modified in Sánchez-Sáez et al. 2021, see their Eq. A5 – to constrain the behaviour in coverage gaps.
@@ 2.3.1 Feature Extraction
X – YRise Time $t_{\texttt{X}\_\texttt{Y}}$ d Time between Xand Yof the peak (e.g. 25 – 75%) 3
Fade Time $t_{Peak\_50}$ d e.g. time between peak and 50% peak flux
Duration Time $t_{50}$ d e.g. time spent above 50% peak flux
@@
Colour $(g-r)_{r,-Xd}$ mag e.g. ZTF $g-r$ calculated X d before r-band peak
@@ 2.4 Galactic and Host Extinction Corrections
We perform a correction for host extinction with the colour at peak, $(g-r)_{g,max}$, for this subset of SNe described above. Relative to V-band ($A_{V}$ = 1 mag) and based on the assumption of a standard Milky Way $R_{V}$ = 3.1 mag (e.g. Cardelli et al., 1989) reddening law (as implemented in python extinction package). We assume an extinction of $A_{g}/A_{V}$ = 1.19 mag, $A_{r}/A_{V}$ = 0.84 mag and $A_{i}/A_{V}$ = 0.61 mag – we ignore the effect of redshift here. For a ZTF $g$ extinction relative to $(g-r)_{g,max}$ reddening of 1 mag, we find an extinction of $A_{g}=3.37$ $(g-r)_{g,max}$ mag. For our correction to ZTF $g$ luminosities, we multiply the $(g-r)_{g,max}$ by 3.37 and apply this to ZTF $g$ magnitudes of all Types. From this point on, we carry through this extinction correction for light curves where $(g-r)_{g,max}$ $>$ 0.25 and $t_{25\_75}$leq 20 d. However, the volumetric-correction weighting applied in Section section 3.2 is based on the uncorrected peak magnitude.
@@ 2.5 Rise Time
We focus on the rise time and how it relates to the CSM as recent studies (e.g. Morozova et al., 2016; Yaron et al., 2017; Morozova et al., 2018; Pearson et al., 2023; Tinyanont et al., 2022; Bruch et al., 2021, 2023; Jacobson-Galán et al., 2023; Hosseinzadeh et al., 2018, 2023; Irani et al., 2024) have shown substantial evidence of the sensitive nature of the rise time to the progenitor properties and pre-explosion conditions (e.g. $M_{CSM}$, $R_{CSM}$, CSM density and progenitor radius).
Traditional rise time measurements – from explosion to peak – require well-constrained explosion epochs through deep non-detections immediately before explosion and good sampling of the early rise. These constraints typically limit the number of events one can reliably measure the explosion epoch of; the early light curve is commonly then fit using a power-law or a polynomial function (Gall et al., 2015; Bruch et al., 2021, 2023). To analyse our heterogeneously sampled dataset, we instead adopt a mathematically defined rise time that does not depend on explosion epoch constraints.
@@ 2.5.1 Rise Distribution
We define rise time ($t_{25\_75}$) as the interval from 25% to 75% of peak flux in ZTF $g$ band. This definition offers several advantages: it’s robust against sampling errors and low S/N measurements; captures the epoch where CSM signatures are strongest; and avoids plateau phases where other physical processes dominate. We select ZTF $g$ band for its sensitivity to CSM interaction signatures in the blue optical (e.g. Groh, 2014; Gal-Yam et al., 2014; Yaron et al., 2017; Kulkarni et al., 2023) – a choice validated by recent modelling of SN 2023ixf, where $g$ band provided optimal fits (minimum $\chi^{2}$ per progenitor model) across all bands (Moriya & Singh, 2024).
Events peaking at 18.5 mag have 25% peak flux at 20th mag (i.e., $+$ 1.5 mag), which does not typically exceed the ZTF detection limit under favourable conditions but may fall below the threshold during sub-optimal conditions (e.g. bright time). Thus, this ensures we capture a substantial portion of the rise whilst remaining sensitive to fainter objects. Fig. 1 shows our measurement methodology.
@@
From our sample of SNe with forced photometry (as of December 31st 2023), we identified 1323 CCSNe that meet the quality criteria established by Perley et al. (2020). Of these, 981 are hydrogen-rich CCSNe (including regular Type II/IIP, IIn, IIb, and H-rich superluminous SNe; SLSNe), while the remaining 342 are classified as stripped-envelope SNe (SESNe). The other SNe in the quality sample, 4009, are Type Ia SNe (see Table 3). Our Type IIb sample is relatively small due to classification challenges inherent to this subtype. The limited spectral resolution of the SEDM makes it difficult to identify the characteristic evolution – specifically, the disappearance of H features and the emergence of often weak He lines in later spectra. Comprehensive classification typically requires multiple spectra obtained at different epochs, which is not always feasible. While this likely results in some incompleteness in our Type IIb sample, the impact on our overall study conclusions is minimal, as these events represent a small fraction of the H-rich SN population.

### v3

@@ Abstract
transients: supernovae – stars: mass-loss
@@ 1 Introduction
Light curves of core-collapse supernovae (CCSNe), Type II SNe in particular, exhibit a large amount of diversity, varying across orders of magnitude in rise times, luminosities, and durations. The simple progenitor scenario, in which the initial mass is the only factor affecting the SN type or its light curve, cannot adequately explain the extensive observational diversity we see in photometry and spectroscopy – particularly with the acknowledgement of the role binarity plays in stellar evolution (e.g., Eldridge et al., 2008; Sana et al., 2012; Eldridge et al., 2018; Zapartas et al., 2019; Zapartas et al., 2021) via binary induced mass transfer and mergers.
An area being explored in greater detail is the degree to which diversity arises from stars with similar initial masses and evolutionary histories that, nonetheless, produce distinct observational signatures at the time of explosion; e.g., varying mass of H envelopes, progenitor radii and H-richness of the outer envelope (e.g., Popov, 1993; Chieffi et al., 2003; Young, 2004; Reynolds et al., 2020; Humphreys et al., 2020; Hiramatsu et al., 2021; Moriya et al., 2023; Dessart & Jacobson-Galán, 2023). Type II SNe result from the core-collapse of stars with initial masses between 8–20 $\textup{M}_{\odot}$ (e.g., Eldridge & Tout, 2004; Smartt, 2009, 2015; Van Dyk, 2017; Beasor et al., 2020). The most common subtype, Type IIP, originate from red supergiants (RSGs) – a connection confirmed through pre-explosion HST imaging (see Smartt, 2009, 2015). Their light curves exhibit H-recombination powered $\sim$100 d plateaus following steep rises to peak brightness, typically occurring within days to a week (e.g., Langer, 2012; Anderson et al., 2014; González-Gaitán et al., 2015; Rubin et al., 2016; Valenti et al., 2016).
@@
From the emergence of narrow emission lines in early spectra of young SNe (flash ionisation; Gal-Yam et al., 2014; Groh, 2014; Yaron et al., 2017; Gal-Yam, 2017; Bruch et al., 2023), strong evidence has been presented for the presence of a substantial mass of dense material close to the progenitor at the time of core-collapse. Narrow lines are likely the result of shock breakout (SBO) shock-heating and ionising a slow-moving, dense material (e.g., Yaron et al., 2017; Irani et al., 2024; Jacobson-Galán et al., 2024a). As the narrow lines typically persist for only a $\sim$ few days post-explosion, it is assumed that the CSM required is the result of mass-loss from the star in the years immediately preceding core-collapse (e.g., Das & Ray, 2017; Davies et al., 2022; Tinyanont et al., 2022; Pearson et al., 2023).
@@
Measurements of the CSM properties from the flash ionisation allow for constraints on the mass-loss rate, $\dot{M}$, and late-stage RSG instabilities experienced in the centuries-decades-years immediately before core-collapse (e.g., Mauron & Josselin, 2011; Yaron et al., 2017; Morozova et al., 2018; Stroh et al., 2021; Pearson et al., 2023; Tinyanont et al., 2022; Bruch et al., 2021, 2023; Moriya et al., 2023). These analyses typically assume that the CSM is an unbound, spherically symmetric material escaping with velocities of order $\sim$10 – 100 km s-1 (Smith, 2014; Morozova et al., 2017), following a density profile that decreases with radius ($\rho\propto r^{-2}$ for steady-state mass loss Morozova et al., 2018; Moriya et al., 2018; Davies et al., 2022; Moriya et al., 2023). Bruch et al. (2021, 2023) find that $\sim$ 60% of Type II SNe show evidence for significant amounts of dense CSM confined to a region around the progenitor at the time of explosion – although, this figure is not corrected for observational biases and not volume limited. Potential precursor events (e.g., Fraser et al., 2013; Jacobson-Galán et al., 2022; Dong et al., 2024; Warwick et al., 2025) provide further evidence of eruptions close to the ‘classical’ core-collapse event.
@@
The notion that many Type IIP SNe progenitors are surrounded by dense CSM at the time of explosion is further supported by detailed studies of nearby events: SN 2021yja ($\approx$23 Mpc; Kozyreva et al., 2022; Hosseinzadeh et al., 2022), SN 2023ixf ($\approx$7 Mpc; Bostroem et al., 2023; Hosseinzadeh et al., 2023; Hiramatsu et al., 2023; Jencson et al., 2023; Jacobson-Galán et al., 2023; Li et al., 2024; Singh et al., 2024; Zimmerman et al., 2024) and SN 2024ggi ($\approx$7 Mpc; Chen et al., 2024, 2025; Xiang et al., 2024; Pessi et al., 2024; Jacobson-Galán et al., 2024b; Shrestha et al., 2024), which, combined with early photometric and spectroscopic data, confirm CSM around their RSG progenitors. In these cases, dense, optically thick CSM causes the SBO to occur within the CSM rather than at the stellar surface (Förster et al., 2018; Tinyanont et al., 2022; Pearson et al., 2023), producing rapid rises and enhanced peak luminosities (e.g., Moriya et al., 2011; Das & Ray, 2017; Morozova et al., 2017, 2018; Bruch et al., 2021, 2023; Tinyanont et al., 2022; Pearson et al., 2023; Moriya et al., 2023; Li et al., 2024).
@@
$\dot{M}$ for RSGs have been measured through multiple techniques: mid-IR observations of circumstellar dust in clusters show $\dot{M}$ $\sim$10-6 – 10-5 $\textup{M}_{\odot}$ yr-1 (e.g., Beasor & Davies, 2018; Beasor et al., 2020), consistent with rates derived from molecular line and radio measurements (e.g., Mauron & Josselin, 2011) and comparing pre-explosion progenitor properties to theoretical stellar evolution models (e.g., Smartt, 2009). Type IIn progenitors exhibit much higher rates of 10-3 – 1 $\textup{M}_{\odot}$ yr-1, derived from multi-wavelength observations (e.g., Kiewe et al., 2012; Taddia et al., 2013; Fransson et al., 2014), and combined X-ray, radio, and spectroscopic signatures (e.g., Smith, 2017a, b).
@@
However, $\dot{M}$ inferred from RSG observations alone are insufficient to produce the measured $M_{CSM}$ and $R_{CSM}$ on the timescale of decades to months pre-explosion (e.g., Bruch et al., 2021, 2023). Popular mechanisms for end-of-life mass-loss include: wave-driven energy heating into the stellar envelope (e.g., Fuller, 2017; Morozova et al., 2020; Wu & Fuller, 2021), radiation-driven mass-loss (e.g., Vink, 2008; Vink, Jorick S. & Sabhahit, Gautham N., 2023), instabilities caused by explosive shell burning (e.g., Arnett & Meakin, 2011; Smith & Arnett, 2014), common envelope interactions caused by binary interactions (e.g., Chevalier, 2012; Sana et al., 2012) and RSG ‘superwinds’ (e.g., Davies et al., 2022).
@@
Previous studies have been successful in characterising smaller subsets of Type II SN properties such as luminosities and rise times (e.g., Taddia et al., 2013; Anderson et al., 2014; Sanders et al., 2015; Gall et al., 2015; Rubin et al., 2016; Valenti et al., 2016; Graur et al., 2017; Davis et al., 2019) but are limited to small numbers or incomplete samples, made up of well-observed SNe detected in heterogeneous galaxy-targeted surveys. Infrequent and inconsistent survey cadence lead to inadequate coverage on the rise, limiting the amount of information one can infer from the rising light curve. The wide area, high cadences and untargeted nature of modern surveys allow for larger, more complete samples to be curated – lending to more detailed statistical analysis of population characteristics and their frequencies. This work builds upon previous studies by making use of robust statistical methods and large, highly complete surveys.
@@ 2.3 Gaussian Process Regression
CCSN light curves are difficult to model due to the extensive variety in their photometric behaviour. There exist analytical attempts to address the problem using parametric fitting functions (e.g., Villar et al., 2017; Villar et al., 2019) and there are advancements in theoretical models to produce synthetic light curves with more likeness to observed light curves (Morozova et al., 2017; Das & Ray, 2017; Moriya et al., 2023, and references therein). Whilst certain parameterisation and generalised empirical models have proven to be adequate in some scenarios, few of these models can fully characterise the diversity of parameters and properties present in transients being uncovered by large surveys. Thus, we are motivated to use a non-parametric technique such as Gaussian process regression (GPR; Rasmussen et al. 2004).
GPR is a non-parametric, Bayesian machine learning method for modelling data with functions of an unknown form (see Aigrain & Foreman-Mackey, 2023, for a review). For single-band SN light curve interpolation, the unknown function a 1-dimensional Gaussian process (GP) approximates is flux as a function of time. We include the effective wavelength, $\lambda_{\rm{eff}}$, of each filter band, and train in 2-dimensions – e.g., flux as a function of time and effective wavelength – which is often expressed as probability by Eq. 1, similar to methodology used in Thornton et al. (2024).
@@
For modelling diverse SN light curves, we use a Matèrn-5/2 covariance function (K) that captures both smooth evolution in addition to the sharp transitions characteristic of SNe – the kernel includes an additive white noise term to account for photometric uncertainties . We implement the GPR using the python package george (Ambikasaran et al., 2015), combining the Matèrn-5/2 kernel with the function from Villar et al. (2019) – modified in Sánchez-Sáez et al. (2021), see their Eq. A5 – to constrain the behaviour in coverage gaps.
@@ 2.3.1 Feature Extraction
X – YRise Time $t_{\texttt{X}\_\texttt{Y}}$ d Time between Xand Yof the peak (e.g., 25 – 75%) 3
Fade Time $t_{Peak\_50}$ d e.g., time between peak and 50% peak flux
Duration Time $t_{50}$ d e.g., time spent above 50% peak flux
@@
Colour $(g-r)_{r,-Xd}$ mag e.g., ZTF $g-r$ calculated X d before r-band peak
@@ 2.4 Galactic and Host Extinction Corrections
We perform a correction for host extinction with the colour at peak, $(g-r)_{g,max}$, for this subset of SNe described above. Relative to V-band ($A_{V}$ = 1 mag) and based on the assumption of a standard Milky Way $R_{V}$ = 3.1 mag (e.g., Cardelli et al., 1989) reddening law (as implemented in python extinction package). We assume an extinction of $A_{g}/A_{V}$ = 1.19 mag, $A_{r}/A_{V}$ = 0.84 mag and $A_{i}/A_{V}$ = 0.61 mag – we ignore the effect of redshift here. For a ZTF $g$ extinction relative to $(g-r)_{g,max}$ reddening of 1 mag, we find an extinction of $A_{g}=3.37$ $(g-r)_{g,max}$ mag. For our correction to ZTF $g$ luminosities, we multiply the $(g-r)_{g,max}$ by 3.37 and apply this to ZTF $g$ magnitudes of all Types. From this point on, we carry through this extinction correction for light curves where $(g-r)_{g,max}$ $>$ 0.25 and $t_{25\_75}$leq 20 d. However, the volumetric-correction weighting applied in Section section 3.2 is based on the uncorrected peak magnitude.
@@ 2.5 Rise Time
We focus on the rise time and how it relates to the CSM as recent studies (e.g., Morozova et al., 2016; Yaron et al., 2017; Morozova et al., 2018; Pearson et al., 2023; Tinyanont et al., 2022; Bruch et al., 2021, 2023; Jacobson-Galán et al., 2023; Hosseinzadeh et al., 2018, 2023; Irani et al., 2024) have shown substantial evidence of the sensitive nature of the rise time to the progenitor properties and pre-explosion conditions (e.g., $M_{CSM}$, $R_{CSM}$, CSM density and progenitor radius).
Traditional rise time measurements – from explosion to peak – require well-constrained explosion epochs through deep non-detections immediately before explosion and good sampling of the early rise. These observational constraints significantly limit the number of events for which explosion epochs can be reliably determined, though this is partially alleviated by fitting the early light curve with power-law or polynomial functions to approximate the explosion time (Gall et al., 2015; González-Gaitán et al., 2015; Bruch et al., 2021, 2023). To analyse our heterogeneously sampled dataset, we instead adopt a mathematically defined rise time that does not depend on explosion epoch constraints.
@@ 2.5.1 Rise Distribution
We define rise time ($t_{25\_75}$) as the interval from 25% to 75% of peak flux in ZTF $g$ band. This definition offers several advantages: it’s robust against sampling errors and low S/N measurements; captures the epoch where CSM signatures are strongest; and avoids plateau phases where other physical processes dominate. We select ZTF $g$ band for its sensitivity to CSM interaction signatures in the blue optical (e.g., Groh, 2014; Gal-Yam et al., 2014; Yaron et al., 2017; Kulkarni et al., 2023) – a choice validated by recent modelling of SN 2023ixf, where $g$ band provided optimal fits (minimum $\chi^{2}$ per progenitor model) across all bands (Moriya & Singh, 2024).
Events peaking at 18.5 mag have 25% peak flux at 20th mag (e.g., $+$ 1.5 mag), which does not typically exceed the ZTF detection limit under favourable conditions but may fall below the threshold during sub-optimal conditions (e.g., bright time). Thus, this ensures we capture a substantial portion of the rise whilst remaining sensitive to fainter objects. Fig. 1 shows our measurement methodology.
@@
From our sample of SNe with forced photometry (as of December 31st 2023), we identified 1323 CCSNe that meet the quality criteria established by Perley et al. (2020). Of these, 981 are hydrogen-rich CCSNe (including regular Type II/IIP, IIn, IIb, and H-rich superluminous SNe; SLSNe), while the remaining 342 are classified as stripped-envelope SNe (SESNe). The other SNe in the quality sample, 4009, are Type Ia SNe (see Table 3). Our Type IIb sample is relatively small due to classification challenges inherent to this subtype. The limited spectral resolution of the SEDM makes it difficult to identify the characteristic evolution – specifically, the disappearance of H features and the emergence of often weak He lines in later spectra. Comprehensive classification typically requires multiple spectra obtained at different epochs, which is not always feasible. While this likely results in some incompleteness in our Type IIb sample, the impact on our overall study conclusions is minimal, as these events represent a small fraction of the H-rich SN population. We discuss and quantify this systematic impact of this in section A.10.